OpenStudy (e_s_j_f):

hey, can someone please help me with this question intregral = x/(3+x^4) dx

OpenStudy (e_s_j_f):

\[\int\limits \frac{ x }{ 3+x^{2} } dx\]

OpenStudy (e_s_j_f):

i am stuck at \[\frac{ 1 }{ 2 } \int\limits \frac{ 1 }{ 3+u^ 2}\]

OpenStudy (e_s_j_f):

forgo du

OpenStudy (asnaseer):

I assume your original question should be:\[\int\limits \frac{ x }{ 3+x^{4} } dx\]

OpenStudy (asnaseer):

and that you then used the substitution:\[u=x^2\]to end up with:\[\frac{ 1 }{ 2 } \int \frac{ 1 }{ 3+u^ 2}du\]

OpenStudy (asnaseer):

your next step should be to factor out the 3 from the denominator to get:\[\frac{1}{6} \int \frac{1}{1+\frac{u^ 2}{3}}du\]

OpenStudy (asnaseer):

finally use the substitution:\[s=\frac{u}{\sqrt{3}}\]

OpenStudy (asnaseer):

hopefully you can complete from here

OpenStudy (e_s_j_f):

why do i have to factor out 3?

OpenStudy (asnaseer):

to try and get it into one of the standard forms

OpenStudy (e_s_j_f):

where did u get u/sqrt(3)

OpenStudy (asnaseer):

I am trying to get it to the form:\[\int\frac{1}{1+x^2}dx=\arctan(x)+C\]

OpenStudy (asnaseer):

\[\frac{u^2}{3}=\left(\frac{u}{\sqrt{3}}\right)^2\]

OpenStudy (asnaseer):

does it make sense now?

OpenStudy (e_s_j_f):

umm kinda, so do i just use 1/3 arctan(x/3) + C

OpenStudy (asnaseer):

use the "s" substitution I gave above. The integral will ten come out in terms of "s". You then need to convert back to "x".

OpenStudy (asnaseer):

*then

OpenStudy (e_s_j_f):

I = 1/sqrt(3) arctan u/sqrt(3) + C

OpenStudy (e_s_j_f):

I = 1/2 (1/sqrt(3) arctan x/sqrt(3)) + C?

OpenStudy (asnaseer):

not quite - can you please show all your steps so that I can spot where you may have made a mistake?

OpenStudy (e_s_j_f):

@asnaseer its x^2 right.

OpenStudy (e_s_j_f):

using trig identiy tan^2x +1 = sec^2x rigjt.

OpenStudy (asnaseer):

are you stating that the final answer to this integral is \(x^2\)?

OpenStudy (e_s_j_f):

I = 1/2 (1/sqrt(3) arctan x^2/sqrt(3)) + C? this one

OpenStudy (asnaseer):

yes :) although a simpler way to write that would be:\[\frac{\arctan(\frac{x^2}{\sqrt{3}})}{2\sqrt{3}}+C\]

OpenStudy (e_s_j_f):

oh i see thanks!

OpenStudy (asnaseer):

yw :)